OpenStudy (anonymous):

can you help me with this one and let me know if I do this correctly? 5x+5y=-11 7x-3y=13 x=12 y=8 5(12)+5(8)=100 7(12)-3(8)=60 what do I need to do?

OpenStudy (anonymous):

were the values of x and y given or did you find them out?

OpenStudy (radar):

You need to review your work. Obviously x=12 and y=8 don't check out!

OpenStudy (anonymous):

I am having trouble with this

OpenStudy (radar):

Were you directed to use "elimination" or substitution to solve this.

OpenStudy (anonymous):

elimination

OpenStudy (radar):

Which one (x or y) do you want to eliminate.....your cholice.

OpenStudy (anonymous):

I am not sure... This is where I guess I am getting confused

OpenStudy (radar):

It doesn't matter, just pick one x or y.

OpenStudy (anonymous):

y

OpenStudy (radar):

Good choice. Now you want to make the coefficient (number) of y in each equation the same. You can multiply in this case, Multiply the first equation (every term) by 3 and in the second equation multiply it by 5. That will make the coefficient of y the same number (15)

OpenStudy (anonymous):

5x+5y=-11 (1) 7x-3y=13 (2) multiply (1) by 7 and (2) by -5 35x + 35y = -77 (3) -35x + 15y = -65 (4) add (3) and (4) to get 50y = - 142 y = -142/50 y = -71/25 substitute y = -71/25 in (1) -71 5x + 5*----- = -11 25 71 5x - ----- = -11 5 multiplying throughout by 5 25x - 71 = -55 -55 + 71 16 x = -------- = ------ 25 25

OpenStudy (radar):

That is the first step in elimination. Your two equations should look like this: 15x + 15y = -33 35x - 15y = 65

OpenStudy (radar):

Do you follow so far?

OpenStudy (radar):

Harkirat has eliminated the x but we can continue with your choice.

OpenStudy (anonymous):

Ok, I am lost

OpenStudy (radar):

Next step is to add those two equations that we just developed. 15x + 15 Y =-33 35X - 15 y= 65 --------------------- Add. 50x = 32 Notice how y got eliminated and we can solve for x.

OpenStudy (radar):

Did you get lost when we multiplied the first equation by 3. Or are you lost in why in the heck we decided to multiply it by 3 in the first place?

OpenStudy (anonymous):

LOL Yeah why multiplied by 3... I think

OpenStudy (radar):

Yes, a good question. We decided we wanted to eliminate y, (your choice and a good one I might add) We are instructed to use the elimination method. To eliminate y we must manipulate the two equations (legally of course!) so that y has the same coefficient. Then we will either add or subtaract y getting zero, or in effect eliminating y. The coefficient of y in the first equation is +5 (5y) the coefficient of y in the second equation is -3 (-3y) We want to make them the same number and we can do that by multiplying the first by 3..........but to keep it legal, we must multiply every term by three or we destroy the balance of the equation. left = right Same goes for the second equation except we multiply thru by 5 Now they are the same coefficient (15) one + and one is a minus. Clear

OpenStudy (anonymous):

ah I think I understand it better now

OpenStudy (radar):

Now we can just add and eliminate y getting 0y 15x +15y=-33 35x - 15y =65 --------------- Adding to eliminate y 50 x +0y =32 50x=32 x=32/50=16/25 same value harkirat got for x.

OpenStudy (radar):

Please review harikart's work and think about this and it will began to make sense.

OpenStudy (anonymous):

so an ordered pair would look like this 32/50,16/25?

OpenStudy (radar):

Let me look at harkirat's work . Harkirats shows a value of y at -71/25. So the ordered pair would be (16/25, -71/25). I haven't doublechecked his work but just lookin it seems o.k.

OpenStudy (radar):

You can verify by substituting for 16/25 for x and solving for y. Hopefully you will come up with the same.

OpenStudy (anonymous):

ok, I am going to try this one 5x+3y=-11 7x-2y=17 I am going to eliminate Y and solve for X (or at least I am going to try)

OpenStudy (radar):

5x+5y=-11 Original equation. 5(16/25) +5y=-11 Substituting 16/25 for x. 80/25 +11 =-5y 80 275 ---- +------ =-5y 25 25 355 ---- = -5y 25 355 -1 y = ---- x ----- =-71/25 25 5

OpenStudy (anonymous):

2(5x + 3y = -11)---->10x + 6y = -22 3(7x - 2y = 17)------>21x - 6y = 51 31x = 29 x = 29/31